OpenStudy (anonymous):

simplified form of (a / 3) / (6 / 3a) simplified form of (a / 3) / (6 / 3a) @Mathematics

OpenStudy (anonymous):

\[\left(\frac{a}{3}\right)/\left(\frac{6}{(3a)}\right)=\frac{a^2}{6} \]

OpenStudy (anonymous):

Thank you